OpenStudy (anonymous):

how do I write an equivalent exponential or logarithmic equation for the following: e^x=55 In44=3.78 e^-3=0.05?

OpenStudy (hunus):

The natural log of a number means e to what power equals this number \[\ln(e^2) = 2\] This says e to what power is equal to e^2 which is obviously 2 \[e^a = b\]is the same thing as \[\ln(b) = a\]

OpenStudy (unklerhaukus):

\[\boxed{\log_bx=y\quad\iff\quad x=b^y}\]\[\boxed{\quad\ln x=y\quad\iff\quad x=e^y}\]